### v1

@@ Abstract
In this paper, we consider a class of structured nonsmooth fractional minimization, where the first part of the objective is the ratio of a nonnegative nonsmooth nonconvex function to a nonnegative nonsmooth convex function, while the second part is the difference of a smooth nonconvex function and a nonsmooth convex function. This model problem has many important applications, for example, the scale-invariant sparse signal recovery in signal processing. However, the existing methods for fractional programs are not suitable for solving this problem due to its special structure. We first present a novel nonfractional min-max reformulation for the original fractional program and show the connections between their global (local) optimal solutions and stationary points. Based on the reformulation, we propose an alternating maximization proximal descent algorithm and show its subsequential convergence towards a critical point of the original fractional program under a mild assumption. By further assuming the Kurdyka-Łojasiewicz (KL) property of an auxiliary function, we also establish the convergence of the entire solution sequence generated by the proposed algorithm. Finally, some numerical experiments on the scale-invariant sparse signal recovery are conducted to demonstrate the efficiency of the proposed method.
keywords:
nonsmooth fractional program, min-max optimization, proximal algorithm, scale-invariant sparse recovery, Kurdyka-Łojasiewicz property
{AMS}
90C32, 90C26, 90C30, 65K05
@@ 1 Introduction
(1) $\min\left\{\frac{f(x)}{g(x)}+h(x):x\in\Omega\cap\mathcal{C}\right\},$
where $f,g:\mathbb{R}^{n}\rightarrow[0,+\infty)$ and $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are proper closed, $\mathcal{C}\subseteq\mathbb{R}^{n}$ is closed and $\mathcal{C}\cap\Omega\neq\emptyset$ with $\Omega:=\left\{x\in\mathbb{R}^{n}:g(x)\neq 0\right\}$. Throughout this paper, we adopt the following blanket assumption for problem (1).
@@
    The proximal operator associated with the sum of $f$ and the indicator function on $\mathcal{C}$ can be evaluated. [^1] The proximal operator associated with the sum of $f$ and the indicator function on $\mathcal{C}$ is defined as $\mathrm{prox}_{f+\iota_{\mathcal{C}}}(x):=\arg\min\{f(z)+\frac{1}{2}\|x-z\|^{2}_{2}:z\in\mathcal{C}\}.$
Problem (1) covers many important optimization models in machine learning and scientific computing. In this work, we focus on its application to the scale-invariant sparse signal recovery [32, 37, 38, 39, 40, 42, 43]. Let $A\in\mathbb{R}^{m\times n}$ be the sensing matrix, $b\in\mathbb{R}^{m}$ be the possibly noisy measurement, $\underline{x}\in\mathbb{R}^{n}$ and $\overline{x}\in\mathbb{R}^{n}$ be the lower bound and upper bound of the underlying signal, respectively. We are particularly interested in the $L_{1}/L_{2}$ robust signal recovery model
(2) $\min\left\{\frac{\|x\|_{1}}{\|x\|_{2}}+\frac{\lambda}{2}\mathrm{dist}^{2}\left(Ax-b,\mathcal{S}_{\mu}\right):x\neq 0,~{}\underline{x}\leq x\leq\overline{x},~{}x\in\mathbb{R}^{n}\right\},$
@@
(3) $\min\left\{\frac{\|x\|_{1}}{\|x\|_{(K)}}+\frac{\lambda}{2}\mathrm{dist}^{2}(Ax-b,\mathcal{S}_{\mu}):x\neq 0,~{}\underline{x}\leq x\leq\overline{x},~{}x\in\mathbb{R}^{n}\right\},$
@@
Now we briefly review two categories of approaches for solving structured nonsmooth fractional optimization problems. The first category of approaches is based on the classical Dinkelbach’s method [14, 17, 19, 35], which is widely used for single-ratio fractional programs. Note that problem (1) can be naively converted into a signal-ratio optimization problem with the numerator $f+hg$, for which the Dinkelbach’s method generate the new iterate $x^{k+1}$ by solving
(4) $\min\left\{f(x)+(hg)(x)-\theta_{k}g(x):x\in\mathcal{C}\right\},$
where $\theta_{k}\in\mathbb{R}$ is updated via $\theta_{k}:=f(x^{k})/g(x^{k})+h(x^{k})$. This iterative scheme is not practical in general since it may be very difficult and expensive to solve an optimization problem of type (4) in each iteration. Recently, proximal algorithms based on the Dinkelbach’s method have been developed for a class of nonsmooth single-ratio fractional programs programs [8, 9, 25, 44]. If we ignore the line-search and extrapolation involved, then applying these algorithms to problem (1) yields the iteration scheme
(5) $x^{k+1}\in\arg\min\left\{f(x)+(hg)(x)-\left\langle\theta_{k}y^{k},x-x^{k}\right\rangle+\frac{\|x-x^{k}\|_{2}^{2}}{2\eta_{k}}:x\in\mathcal{C}\right\},$
@@
(6) $x^{k+1}\in\arg\min\left\{f(x)+\left\langle\nabla(hg)(x^{k})-\theta_{k}y^{k},x-x^{k}\right\rangle+\frac{\|x-x^{k}\|_{2}^{2}}{2\eta_{k}}:x\in\mathcal{C}\right\},$
@@
where $y^{k}\in\partial g(x^{k})$, $\eta_{k}>0$ is some stepsize, and $\theta_{k}$ is defined as in (4). It is worth noting that (5) actually requires that the proximal operator associated with the sum of $f+hg$ and indicator function on $\mathcal{C}$ can be evaluated, while (6) assumes the smoothness of $hg$. Unfortunately, the above requirements do not hold in general and consequently the iteration schemes (5) and (6) are not suitable for tackling problem (1). Therefore, we conclude that the Dinkelbach’s method and its existing variants are not applicable to solving problem (1). The second category of approaches are based on the so-called quadratic transform [5, 10, 36]. This transform introduces an additional variable and provides an equivalent nonfractional maximization problem for fractional programs that maximizes the sum of several ratios. Specifically, it can be applied to the alternative problem that is obtained via replacing the minimum in problem (1) by the maximum, that is,
(7) $\max\left\{\frac{f(x)}{g(x)}+h(x):x\in\Omega\cap\mathcal{C}\right\}.$
Consequently, the equivalent nonfractional problem for (7) is formulated as
(8) $\max\left\{2y\sqrt{f(x)}-y^{2}g(x)+h(x):x\in\Omega\cap\mathcal{C},~{}y\in\mathbb{R}\right\}.$
Since for a fixed $x\in\Omega\cap\mathcal{C}$ the optimal objective of (8) is attained at $y=\sqrt{f(x)}/g(x)$, one can easily utilize block coordinate ascent methods for efficiently solving (8). However, the minimization problem (1) is generally independent of the maximization problem (7) due to the non-negativity requirements of their numerators and denominators. Hence, the quadratic transform can not be directly used to construct an equivalent nonfractional minimization problem for (1).
@@
We also note that some algorithms have been specifically designed for solving the scale-variant sparse signal recovery model, which are based on neither the Dinkelbach’s method nor the quadratic transform. The alternating direction method of multipliers (ADMM) has been applied to the noiseless $L_{1}/L_{2}$ signal recovery model and the model (2) with $\mu=0$ via various variable-splitting schemes [32, 39]. However, these ADMM-type algorithms rely on the structure of the $L_{1}/L_{2}$ function and can not be extended to deal with problem (1) which admits a more general setting for $f$ and $g$. For example, they can not be generalized to solve model (3) due to the nonsmoothness of $\|\cdot\|_{(K)}$. Very recently, the gradient descent flow algorithm (GDFA) is proposed for models (2) and (3) with $\mu=0$ [38]. Although the efficiency of GDFA is demonstrated in the numerical experiments of [38], so far its convergence remains unknown to the best of our knowledge.
@@
Motivated by the above analysis, we aim at designing efficient numerical methods with guaranteed convergence for solving problem (1) by exploiting its intrinsic structure. Specifically, the contributions of this paper are summarized as follows. First, we present an novel nonfractional min-max reformulation for problem (1), which is in the form of
[…]
This reformulation is equivalent to (1) in the sense that $x^{\star}\in\mathbb{R}^{n}$ is a global (local) minimizer of problem (1) if and only if $(x^{\star},c_{\star})$ is a global (local) min-max point of problem (9) for some $c_{\star}\in\mathbb{R}$. Moreover, we prove that $x^{\star}$ is a (Fréchet) stationary point of problem (1) if and only if $(x^{\star},c_{\star})$ is a stationary point of problem (9) for some $c_{\star}\in\mathbb{R}$. To the best of our knowledge, problem (9) is the first equivalent min-max reformulation for fractional programs of the type (1). Second, we propose an alternating maximization proximal descent algorithm (AMPDA) for solving the min-max problem (9). Problem (9) falls into the class of nonsmooth nonconvex-concave min-max problems, which has received considerable attention recently; see, for example, [7, 16, 18, 22, 24, 26] and the references therein. However, existing min-max optimization methods for this class are not suitable for solving problem (9) since for fixed maximization variable $c\in\mathbb{R}$ its objective function involves a nonsmooth nonconvex product term $-c^{2}f(x)g(x)$ with respect to the minimization variable $x$. Hence, our development of the AMPDA is nontrivial. In particular, we ingeniously employ the majorization-minimization scheme and the line-search technique to perform a proximal update associated with the sum of $f$ and the indicator function on $\mathcal{C}$ for the variable $x$, where the line-search procedure determines the proximal stepsize and is terminated when some sufficient descent condition is satisfied. We show that any accumulation point of the solution sequence generated by the AMPDA is a critical point of problem (1) under a mild assumption. Lastly, by further assuming that some auxiliary function is a proper closed KL function, we prove the convergence of the entire solution sequence generated by AMPDA. It is worth noting that our sequential convergence analysis for AMPDA does not require any additional assumption on the denominator $g$, while those for nonsmooth fractional optimization methods in existing literature always assume that either the conjugate of $g$ is continuous on its domain or $g$ is locally Lipschitz differentiable on $\Omega$; see, for example, [9, 25].
@@
The rest of this paper is organized as follows. In Section 2, we introduce notation and present some preliminary materials. In Section 3, we study the relationships between problem (1) and its min-max reformulation (9). We propose the AMPDA and establish its subsequential convergence in Section 4. The sequential convergence of the AMPDA is further investigated in Section 5. Finally, we present some numerical results of the AMPDA for models (2) and (3) in Section 6.

### v2

@@ Abstract
In this paper, we consider a class of structured nonsmooth fractional minimization, where the first part of the objective is the ratio of a nonnegative nonsmooth nonconvex function to a nonnegative nonsmooth convex function, while the second part is the difference of a smooth nonconvex function and a nonsmooth convex function. This model problem has many important applications, for example, the scale-invariant sparse signal recovery in signal processing. However, the existing methods for fractional programs are not suitable for solving this problem due to its special structure. We first present a novel nonfractional min-max reformulation for the original fractional program and show the connections between their global (local) optimal solutions and stationary points. Based on the reformulation, we propose an alternating maximization proximal descent algorithm and show its subsequential convergence towards a critical point of the original fractional program under a mild assumption. Moreover, we prove that the proposed algorithm can find an $\epsilon$-critical point of the considered problem within $\mathcal{O}(\epsilon^{-2})$ iterations. By further assuming the Kurdyka-Łojasiewicz (KL) property of an auxiliary function, we also establish the convergence of the entire solution sequence generated by the proposed algorithm. Finally, some numerical experiments on the $L_{1}/L_{2}$ least squares problem and scale-invariant sparse signal recovery are conducted to demonstrate the efficiency of the proposed method.
@@ 1 Introduction
$\min\left\{\frac{f(x)}{g(x)}+h(x):x\in\Omega\cap\mathcal{C}\right\},$ (1)
where $f,g:\mathbb{R}^{n}\rightarrow[0,+\infty)$ and $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ are proper closed, $\mathcal{C}\subseteq\mathbb{R}^{n}$ is closed and $\mathcal{C}\cap\Omega\neq\emptyset$ with $\Omega:=\left\{x\in\mathbb{R}^{n}:g(x)\neq 0\right\}$. Let $f_{\mathcal{C}}:\mathbb{R}^{n}\to[0,+\infty]$ be the sum of $f$ and the indicator function on $\mathcal{C}$. Throughout this paper, we adopt the following blanket assumption for problem (1).
@@
    The proximal operator associated with $f_{\mathcal{C}}$ can be evaluated. [^1] The proximal operator associated with $f_{\mathcal{C}}$ is defined as $\mathrm{prox}_{f_{\mathcal{C}}}(x):=\arg\min\{f_{\mathcal{C}}(z)+\frac{1}{2}\|x-z\|^{2}_{2}:z\in\mathbb{R}^{n}\}.$
Problem (1) covers many important optimization models in machine learning and scientific computing. In this work, we focus on its application to the scale-invariant sparse signal recovery Rahimi-Wang-Dong-Lou:2019SIAM-SC ; TaoMin:2022SIAM-SC ; Wang-Aujol-Gilboa-Lou:2024IPI ; Wang-Tao-Nagy-Lou:2021SIAM-ImageScience ; Wang-Yan-Rahimi-Lou:2020IEEE ; Yin-Esser-Xin:2014CIS ; Zeng-Yu-Pong:2021SIAM-OPT ; Zhang-Liu-Li:2025ACHA ; NaZhang-QiaLi:2022SIAM-OPT . Let $A\in\mathbb{R}^{m\times n}$ be the sensing matrix, $b\in\mathbb{R}^{m}$ be the possibly noisy measurement, $\underline{x}\in\mathbb{R}^{n}$ and $\overline{x}\in\mathbb{R}^{n}$ be the lower bound and upper bound of the underlying signal, respectively. We are particularly interested in the $L_{1}/L_{2}$ robust signal recovery model
$\min\left\{\frac{\|x\|_{1}}{\|x\|_{2}}+\frac{\lambda}{2}\mathrm{dist}^{2}\left(Ax-b,\mathcal{S}_{\mu}\right):x\neq 0,~\underline{x}\leq x\leq\overline{x},~x\in\mathbb{R}^{n}\right\},$ (2)
@@
$\min\left\{\frac{\|x\|_{1}}{\|x\|_{(K)}}+\frac{\lambda}{2}\mathrm{dist}^{2}(Ax-b,\mathcal{S}_{\mu}):x\neq 0,~\underline{x}\leq x\leq\overline{x},~x\in\mathbb{R}^{n}\right\},$ (3)
@@
Now we briefly review two categories of approaches for solving structured nonsmooth fractional optimization problems as follows.
The first category of approaches is based on the classical Dinkelbach’s method Crouzeix-Ferland-Schaible:1985JOTA ; Dinkelbach-Werner:MS1967 ; Ibaraki:1983Mathematical_Programming ; Schaible:1976Fractional_Programming , which is widely used for single-ratio fractional programs. Note that problem (1) can be naively converted into a signal-ratio optimization problem with the numerator $f+hg$, for which the Dinkelbach’s method generates the new iterate $x^{k+1}$ by solving
$\min\left\{f(x)+(hg)(x)-\theta_{k}g(x):x\in\mathcal{C}\right\},$ (4)
where $\theta_{k}\in\mathbb{R}$ is updated via $\theta_{k}:=f(x^{k})/g(x^{k})+h(x^{k})$. This iterative scheme is not practical in general since it may be very difficult and expensive to solve an optimization problem of type (4) in each iteration. Recently, proximal algorithms based on the Dinkelbach’s method have been developed for a class of nonsmooth single-ratio fractional programs Bot-Csetnek:2017Optimization ; Bot-Dao-Li:2021MathematicsofOperationsResearch ; Li-Shen-Zhang-Zhou:2022ACHA ; NaZhang-QiaLi:2022SIAM-OPT ; Zhang-Liu-Li:2025ACHA ; Zhou-Zhang-Li:2024MOR . If we ignore the line-search and extrapolation involved, then applying these algorithms to problem (1) yields the iteration scheme
$x^{k+1}\in\arg\min\left\{f(x)+(hg)(x)-\left\langle\theta_{k}y^{k},x-x^{k}\right\rangle+\frac{\|x-x^{k}\|_{2}^{2}}{2\eta_{k}}:x\in\mathcal{C}\right\},$ (5)
@@
$x^{k+1}\in\arg\min\left\{f(x)+\left\langle\nabla(hg)(x^{k})-\theta_{k}y^{k},x-x^{k}\right\rangle+\frac{\|x-x^{k}\|_{2}^{2}}{2\eta_{k}}:x\in\mathcal{C}\right\},$ (6)
@@
where $y^{k}\in\partial g(x^{k})$, $\eta_{k}>0$ is some stepsize, and $\theta_{k}$ is defined as in (4). It is worth noting that (5) actually requires that the proximal operator associated with the sum of $f+hg$ and indicator function on $\mathcal{C}$ can be evaluated, while (6) assumes the smoothness of $hg$. Unfortunately, the above requirements do not hold in general and consequently the iteration schemes (5) and (6) are not suitable for tackling problem (1). Therefore, we conclude that the Dinkelbach’s method and its existing variants are not applicable to solving problem (1).
The second category of approaches is based on the so-called quadratic transform Benson:2004JOTA ; Bot-Dao-Li:2023SIAM-OPT ; Shen-Yu:2018IEEE . Let $\varphi_{i},\psi_{i}:\mathbb{R}^{n}\to[0,+\infty)$, $i=1,2,\ldots,N$, be proper closed functions. The quadratic transform refers to reformulating the problem of maximizing a sum of ratios $\sum_{i=1}^{N}\frac{\varphi_{i}(x)}{\psi_{i}(x)}$ into the one of maximizing problem $\sum_{i=1}^{N}2y_{i}\sqrt{\varphi_{i}(x)}-y^{2}_{i}\psi_{i}(x)$ with respect to $y\in\mathbb{R}^{N}$ and $x$. The equivalence of global optimality is built on the fact that $\max_{y_{i}}2y_{i}\sqrt{\varphi_{i}(x)}-y^{2}_{i}\psi_{i}(x)=\frac{\varphi_{i}(x)}{\psi_{i}(x)}$, whereas the equivalence of the stationarity is further investigated in Bot-Dao-Li:2023SIAM-OPT . This transform can be trivially adapted to the following fractional minimization problem
$\min\left\{-\frac{f(x)}{g(x)}+h(x):x\in\Omega\cap\mathcal{C}\right\}.$ (7)
The resulting minimization problem is in the form of
$\min\left\{-2y\sqrt{f(x)}+y^{2}g(x)+h(x):x\in\Omega\cap\mathcal{C},~y\in\mathbb{R}\right\}.$ (8)
and the equivalence between (7) and (8) at the levels of global optimality and stationarity follows immediately. However, the main fractional minimization problem (1) considered in this paper is generally independent of (7) due to the non-negativity requirements imposed on $f$ and $g$. Although one may further reformulate (1) into
$\min_{x\in\Omega\cap\mathcal{C}}\;\max_{y\in\mathbb{R}}\;2y\sqrt{f(x)}-y^{2}g(x)+h(x),$ (9)
by exploiting again the identity $f(x)/g(x)=\max_{y\in\mathbb{R}}2y\sqrt{f(x)}-y^{2}g(x)$, we emphasize that this formulation goes beyond the scope of the quadratic transform, as it is a min-max problem. The equivalence between the stationarity of (1) and that of (9) is in general unknown and requires further investigation. In addition, as in the treatment of problem (8), addressing the min-max problem (9) typically requires $\sqrt{f}$ to be locally Lipschitz continuous, or even convex in some settings. These requirements are substantially stronger than assuming local Lipschitz continuity of $f$ itself, which we adopt for the problem (1). For example, the function $f(x)=\|x\|_{1}$ is convex and locally Lipschitz continuous, whereas $\sqrt{f(x)}=\sqrt{\|x\|_{1}}$ is neither convex nor locally Lipschitz continuous.
@@
We also note that some algorithms have been specifically designed for solving the scale-variant sparse signal recovery model, which are based on neither the Dinkelbach’s method nor the quadratic transform. The alternating direction method of multipliers (ADMM) has been applied to the noiseless $L_{1}/L_{2}$ signal recovery model and the model (2) with $\mu=0$ via various variable-splitting schemes Rahimi-Wang-Dong-Lou:2019SIAM-SC ; Wang-Tao-Nagy-Lou:2021SIAM-ImageScience . However, these ADMM-type algorithms rely on the structure of the $L_{1}/L_{2}$ function and can not be extended to deal with problem (1) which admits a more general setting for $f$ and $g$. For example, they can not be generalized to solve model (3) due to the nonsmoothness of $\|\cdot\|_{(K)}$. Very recently, the gradient descent flow algorithm (GDFA) is proposed for models (2) and (3) with $\mu=0$ Wang-Aujol-Gilboa-Lou:2024IPI . Although the efficiency of GDFA is demonstrated in the numerical experiments of Wang-Aujol-Gilboa-Lou:2024IPI , so far its convergence remains unknown to the best of our knowledge.
@@
Motivated by the above analysis, we aim at designing efficient numerical methods with guaranteed convergence for solving problem (1) by exploiting its intrinsic structure. Specifically, the contributions of this paper are summarized as follows.
First, we present a novel nonfractional min-max reformulation for problem (1)
$\min_{x\in\mathbb{R}^{n}}\max_{c\in\mathbb{R}}\;\widetilde{F}(x,c):=2cf(x)-c^{2}f(x)g(x)+h(x)+\iota_{\Omega\cap\mathcal{C}}(x),$ (10)
where $\iota_{\Omega\cap\mathcal{C}}:\mathbb{R}^{n}\to\{0,+\infty\}$ is the indicator function on $\Omega\cap\mathcal{C}$. The reformulation (10) is equivalent to (1) in the sense that $x^{\star}\in\mathbb{R}^{n}$ is a global (local) minimizer of problem (1) if and only if $(x^{\star},c_{\star})$ is a global (local) min-max point of problem (10) for some $c_{\star}\in\mathbb{R}$. Moreover, we prove that $x^{\star}$ is a (Fréchet) stationary point of problem (1) if and only if $(x^{\star},c_{\star})$ is a stationary point of problem (10) for some $c_{\star}\in\mathbb{R}$. To the best of our knowledge, problem (10) is the first equivalent min-max reformulation for fractional programs of the type (1).
Second, we propose an alternating maximization proximal descent algorithm (AMPDA) for solving the min-max problem (10). Motivated by the min-max reformulation (10), AMPDA adopts a natural alternating strategy: at the $k$-th iteration, fixing $x^{k}$, one maximizes $\widetilde{F}(x^{k},c)$ with respect to $c$, which admits a closed-form update $c_{k}=1/g(x^{k})$; fixing $c_{k}$, one then minimizes the possibly nonsmooth and nonconvex function $\widetilde{F}(x,c_{k})$ with respect to $x$. To handle the $x$-subproblem, we construct a majorization surrogate of $\widetilde{F}(x,c_{k})$ in a neighborhood of the current iterate $x^{k}$. Based on this surrogate, the $x$-update at the $k$-th iteration can be efficiently computed via the proximity operator associated with $f_{\mathcal{C}}$. For many practical applications, including the signal recovery models (2) and (3), this proximity operator can be evaluated efficiently. The proximal stepsize is determined by a line-search procedure, whose stopping criterion is stronger than the standard sufficient descent condition for the objective in (1), but is crucial for establishing the convergence of the entire solution sequence generated by the proposed algorithm. We emphasize that, despite the recent advances in nonsmooth min-max optimization (see, e.g., Bian-Chen:2024MP ; Dai-Wang-Zhang:2024SIAM-OPT ; He-Zhang-Xu:2024JGO ; Kong-Monteiro:2021SIAM-OPT ; Li-Zhu-So:2022NeurIPS-OPT ; Lu-Tsaknakis-Hong-Chen:2020TSP ), existing methods are not suitable for solving the min-max problem (10). The main difficulty arises from the fact that, for a fixed maximization variable $c\in\mathbb{R}$, the objective function involves a nonsmooth and nonconvex product term $c^{2}f(x)g(x)$ with respect to the minimization variable $x$. The construction of the surrogate function in the $x$-update of the proposed AMPDA is therefore not generic but carefully tailored to the specific structure of the term $c^{2}f(x)g(x)$ in the objective of (10), which makes the development of AMPDA highly nontrivial.
Third, we establish the convergence analysis of the proposed AMPDA. We show that any accumulation point of the solution sequence generated by AMPDA is a critical point of problem (1) under a mild assumption. Moreover, we prove that AMPDA can find an $\epsilon$-critical point of problem (1) within $\mathcal{O}(\epsilon^{-2})$ iterations. By further assuming that an auxiliary function is a proper closed KL function, we establish the convergence of the entire solution sequence generated by AMPDA. It is worth noting that our whole-sequence convergence analysis for AMPDA does not require any additional assumption on the denominator $g$, whereas existing nonsmooth fractional optimization methods typically assume that either the conjugate of $g$ is continuous on its domain or $g$ is locally Lipschitz differentiable on $\Omega$; see, for example, Bot-Dao-Li:2021MathematicsofOperationsResearch ; Li-Shen-Zhang-Zhou:2022ACHA .
@@
The rest of this paper is organized as follows. In Section 2, we introduce notation and present some preliminary materials. In Section 3, we study the relationships between problem (1) and its min-max reformulation (10). We propose AMPDA and establish its subsequential convergence in Section 4. The sequential convergence of AMPDA is further investigated in Section 5. Finally, we present some numerical results of AMPDA in Section 6.